OpenStudy (anonymous):

I am currently working with triple integrals. But I have problems figuring out the bounds of the variables. Example question: The solid enclosed by the cylinder x = y^2 and the planes z = 0 and x + z = 1 Evaluate this triple integral

OpenStudy (anonymous):

hmm, x = y^2 is not a cylinder. It is a parabola. And the solid you get bycutting it with the planes is not closed. So volume would be infinite

OpenStudy (anonymous):

My text book gives an answer though, 8/15. I cant figure out the y and the left over x bound.

OpenStudy (anonymous):

The bounds for z are z = 1 - x and z = 0. So the only thing left is the "double" integral for the xy-plane if you take z as a vertical axis. But I can't figure out how to get these left over x and y bounds.

OpenStudy (anonymous):

and that's the problem, because there are no bounds: